Presentation is loading. Please wait.

Presentation is loading. Please wait.

Log for RUN11 1. pedestal threshold scan Data for threshold scan were taken by Nihashi- kun. data: 5data points for theshold 5, 10, 15, 20, 25mV 2.

Similar presentations


Presentation on theme: "Log for RUN11 1. pedestal threshold scan Data for threshold scan were taken by Nihashi- kun. data: 5data points for theshold 5, 10, 15, 20, 25mV 2."— Presentation transcript:

1 log for RUN11 1

2 pedestal threshold scan Data for threshold scan were taken by Nihashi- kun. data: 5data points for theshold 5, 10, 15, 20, 25mV 2

3 pedestal :: comparison with RUN10 almost channels are within 5% difference comparing with RUN10 3ch are out of this range – ch11: 7% up – ch27: 8% down – ch92: 12% up – see detatils Ratio of pedestal : RUN11/RUN10 pedestal :: RUN11, RUN10 In RUN10, we did not exchange any FEM boards and daughter cards, but only exchange fuse on board#10(ch8-15). There is no relation /01/16

4 details 4 ch0 as reference ch11 : 7% up I just mention that the fluctuation of the Adc distribution dose not come from statistics. I think it due to systematics, ex. resolution of internal charge injector or ADC. pedestal = 944.9[ch] pedestal = 707.5[ch] ADC [ch] Q-scan DAC vs ADC Q-scan DAC vs ADC

5 details 5 ch27 : 8% down ch92: 12% up pedestal = 829.5[ch] pedestal = 779.4[ch] The results seems fine. we will use these pedestal for RUN11. ADC [ch] Q-scan DAC vs ADC Q-scan DAC vs ADC

6 trouble?? ch27 has some strange behavior. – like lack of bit keep on eyes. 6 Q-scan taken on Jan 7th laser #324219

7 online monitor modified code (BbcMonDrow.C & BbcMon.C ) to ignore the laser hit rate of 0 for ch68, ch75, ch96, ch97 7 PRDF for test located on /common/s1/TestPRDFF/Run9pp500/EVENTDATA_P PRDFF error message will not be shown even if the laser hit rate is 0 for ch68, 75, 96, 97 (North) 01/21

8 comparison of overflow1 between RUN10 & RUN11 8 typically lower than RUN10 this difference never affect BBLL1 overflow : RUN11 – Run10 [ch] in makebbmap.C timing threshold [ns] for BBLL1 tmax[loop] = (pedestal[loop] - STEP)*tvc2ns[loop] - tzeroAdc[loop]; pedestal -> overflow1 STEP -> 50[ch] tvc2ns -> conversion factor ch to ns

9 to do for time in make pre-LUT including new overflow1 – done. – /data2/phnxbb/USERS/phnxbb/Run11/BBLL1/Ma kingTool/wrk/trig – checked by CheckingTool and seems fine 9

10 HV gain curve scan – -100, -50, 0, +25, +50, +75V 1) lack of laser – ch68, 75, 96, 97 – (ID:: N5, N12, N33, N34) 2) +75 HV point fail – ch29 (ID:: S30) for 1) & 2), used gain curve taken in RUN11 10

11 Trip of 1 HV Group (N-2) 2011/01/23 ~0:00 HV_BB_N-2 – N13,N18,N21,N24,N28, N31, N32 – (ch76, 81, 84, 87, 91, 94, 95) located in outer ring status 40 Trip for voltage error could NOT recover it. need to access 11

12 StationBoard IDLocationCable ID 0NonUpN1 1BBC1UpN3 1BBC1DownN4 2BBC3UpN5 2BBC3DownN6 3NonUpN7 3NonDownN8 4BBC5UpS1 4BBC5DownS2 5BBC6UpS3 5BBC6DownS4 6BBC7UpS5 6BBC7DownS6 7BBC8UpS7 7BBC8DownS8 11BBC2 Down UP N /01/26 IR access : switch BBC2(N-2) to station11-up

13 First collision collision is provided in 1/24 early morning. something wrong?? 13 BBC LL1 delay 14 13

14 issue for LL1 delay All figures show correlation of hit timing of BBC north and south. The difference of left and right is only selected trigger, BBCLL1 or ZDCLL1. We cant simultaneously see collision event in BBC by BBCLL1 and ZDCLL1. 14 BBC LL1 delay = 13 : # : fine delay of v124 = 170 BBC LL1 delay = 14 : # : fine delay of v124 = 170 ArmHitTime North [ns] ArmHitTime South [ns] triggered by BBCLL1 triggered y BBCLL1 triggered by ZDCLL1 collision events! at LL1 delay =13 In contrast, cant see collision events at LL1 delay =14 but not here can see collision events

15 Time in BBCLL1 and ZDCLL1 seem to be timed in correctly 15 BBC LL1 delay = 14 : # : fine delay of v124 = 170 : GL1 Output delay = 5 ArmHitTime North [ns] ArmHitTime South [ns] triggered by BBCLL1 triggered by ZDCLL1 01/25 ~2AM

16 lost MIP peak on ch40 in monhist ch40(S41) : low gain ?? Same HV in RUN10 is applied. : (HV-BB-S7: [V] ) 16 In RUN11 Whats happened!? comparison with ch58(S59) : same HV group, opposite location In RUN10 In RUN10 In RUN11 ch40 ch58

17 raw Adc comparison with RUN10 The shape of the BG is not same between RUN10 & RUN11, but I think its ok since collision type is different. We already know pC/MIP is fluctuate within 20%, But this case, we cant see any MIP peak in RUN11. There is no difference of ch/DAC value between RUN10 and RUN11 taken by Q- scan. (taken in today). And laser gain curve looks fine. I hope to take laser data without beam condition and compare with data taken on Dec. I guess there are difference of peak position of laser in Adc distribution between before and after beam in... In addition, try to apply +50V higher HV. It corresponds to x1.2 higher gain. 17 Cant see MIP peak in RUN11 red:: RUN11 blue:: RUN10 red:: RUN11 blue:: RUN10 ch58(S59) same HV group with ch40 ch40(S41)

18 Adc distribution in collision data The problem is that we cant see MIP peak in ch40 at RUN11. – the Adc distribution at RUN10 is overwritten to compare. below looks fine – Raw signal of laser and thermal noise by oscilloscope – Q-scan and T-scan outputs and gain curve – We already confirmed that PMT gain is almost same with RUN Cant see MIP peak in RUN11 red:: RUN11 blue:: RUN10 red:: RUN11 blue:: RUN10 ch58(S59) same HV group with ch40 ch40(S41) Adc [ch]

19 Adc distribution for Laser (not collision data) 19 ch40 ch58 (as reference) Adc distributions for laser signal are shown for ch40 and ch58. overwritten Adc distribution for 3 runs which setting of HV and laser intensity was not changed. The peaks on ch58 and other channels(except ch40) are stable. In contrast, the peaks for ch40 was often shifted lower side(red). After Low Voltage recycle and compete feed, the peak position come back base position. Adc [ch] NOT working fine… black : base red : next run of base. not changed any setting with base. blue : after LV recycle and complete feed. Adc distribution of Laser signal Adc distribution of Laser signal

20 Tdc0 & Tdc1 timing for Laser (not collision data) Tdc0 & Tdc1 distributions for laser signal are shown for ch40 and ch58. The peaks of laser timing for ch40 was also often shifted(red). We doubt that FEM for ch40 have something trouble. 20 ch40 Tdc0 ch40 Tdc1 ch58 Tdc0 ch58 Tdc1 Tdc [ch] NOT working fine…

21 summary & question The problem is that we cant the MIP peak on Adc distribution of ch40 in collision data. Laser data for Adc, Tdc0 and Tdc1 of ch40 are often and suddenly shifted. – it was fixed after LV recycle and complete feed. We doubt that FEM is NOT working fine. And It lead the problem that we cant see MIP peak. We have IR access in Wednesday morning. – still not find exact origin of the problem. – What can we do to fix this problem at access?? 21

22 Laser signal (in collision data) Adc and Tdc1 distribution(ch40) of Laser taken by PPG(laser) trigger in collision data are shown. expected peak position is known, but cant see on the position. I guess the data of ch40 was always shifted at taking collision data. And it lead the problem that we cant see MIP peak. FEM have something problem. 22 ch40 Tdc1 ch40 Tdc1 expected peak position ~1100ch expected peak position ~2200ch

23 threshold is working ?? in # Black :: threshold=15mV (tdc0) Red :: threshold =30mV (tdc1) pedestal = 802[ch] Looks fine.

24 100V higher HV setting for S-7 We can see correlation of HV value and gain. 24 BBCLL1, |Zvertex|<30, nHitPmt[0]<4, Tdc0<2500 laser event (ScaledTrig&0x ) Blue : Normal HV setting Red : +100V higher

25 IR access : 2011/02/03 : issue(2) 2) replaced ADC daughter card of ch27(S28) – FEM board ID #12, ch3 25 after replacement before replacement ADC [ch] after replacement, ch27 looks fine for Q,T-scan and Laser event. Q-scan

26 IR access : 2011/02/03 : issue(2) change pedestal value. need to pedestal scan again for this ch. 26

27 IR access : 2011/02/03 we exchanged input register of ch40 27 After IR access. (No bema condition), we took laser, T&Q- scan data and it looks fine. But at the collision data taking, the gain of ch40 dropped… I mean problem is still on (cant see MIP peak).

28 gain for ch40 28 Time (2/6) Adc – pedestal [ch] : Laser signal beam on This plot shows the mean position of laser signal for ch40 and ch43 as a function of time Gain of ch40 is unstable. – the gain is drop off when beam is injected. – In addition, gain is getting up after beam lost. 15% up for 30minites beam on 2011/02/06 ch43 ch40 30 minutes

29 gain at collision we found gain of ch40 is extremely and suddenly drop off when both blue and yellow beam is injected. Plots show ADC distribution of laser signal continuously taking data during beam injection. 29 ch40 ADC (Laser signal) pedestal 2011/02/06 growing up Adc[ch] taking laser data continuously both peaks are signal of laser.

30 summary We found the problem appear at collision data. But we cant decouple the source of problem, whether PMT or FEM. If the problem is due to FEM, we have possibilities to be able to fix at next IR access Can you explain this phenomenon by FEM problem?? 30

31 modify injection rate of internal charge injector /export/software/oncs/online_configuration/GTM/ – GTM.BB.1kHz_INITQFA START: # Start of schedular code C;BRANCH # init once, Ax78 # then do nothing for 78x120x106ns, – injected 1./(78 x 120clock x 106ns) ~ 1000Hz B;RESET # then fire test pulse and return to A – need to compile to make GTM.BB.1kHz_INITQFA.gtm modified BBC_Qscan.csh BBC_Qscan_1kHz.csh – # GTM setup – set gtmfile = GTM.BB.1kHz_INITQFA.gtm – set gtmsect = "GTM.BB" 31

32 Q-scan output There is no difference between – Red: 1KHz Q-scan with beam (collision rate : order ~10KHz) – Blue: 1kHz Q-scan without beam – Black: 78Hz Q-scan without beam 32 Adc[ch] 2011/02/08

33 plan for IR access (2/9) only 2 hours 1) measure hit rate dependence by using pulse generator to check FEM 2) re-connect signal cable of ch40 to unused channel of ZDC FEM We did not work in IR. We will do test of cable swap after end of p+p run. High hit rate test will be done at next IR access (2/16) decide to do mask of BBLL1 for ch /02/08

34 34 HV for +- field (30, 40pC/MIP HV) HV_BB_S HV_BB_S HV_BB_S HV_BB_S HV_BB_S HV_BB_S HV_BB_S HV_BB_S HV_BB_N HV_BB_N HV_BB_N HV_BB_N HV_BB_N HV_BB_N HV_BB_N HV_BB_N The operational HV was calculated by using pC/MIP obtained from # S3,S7,N3,N7(inner) are applied 30pC/MIP HV, others are applied 40pC/MIP HV. Valid 2011/02/07 04:35~, #329939~

35 keep in eyes for MIP of these channel # : new HV setting 35 ch25 S26 HV-S3 ch51 S52 HV-S7 ch58 S59 HV-S7 ch83 N20 HV-N3 ch90 N27 HV-N3 ch115 N52 HV-N7

36 MIP for ch40 Red : +100V higher Blue : +150V higher OK, we can see MIP peak in ch40 36

37 SubSystemControl.pl – /export/software/oncs/R11.0.6/install/Linux.i sl5.3.0/bin /home/phoncs/hv_stuff/hvcontrol_BB – quick_init HV 37

38 38 Operational HV RUN11 for +- field (30, 40pC/MIP HV) HV_BB_S HV_BB_S HV_BB_S HV_BB_S HV_BB_S HV_BB_S HV_BB_S HV_BB_S HV_BB_N HV_BB_N HV_BB_N HV_BB_N HV_BB_N HV_BB_N HV_BB_N HV_BB_N The operational HV was calculated by using pC/MIP obtained from # S3,S7,N3,N7(inner) are applied 30pC/MIP HV, others are applied 40pC/MIP HV. Valid 2011/02/10 04:45~, # (calibration)

39 comparing ch/MIP measured ch/MIP value is obtained from # within 10% difference between expected and measured 39 Red: measured #330557, Blue : expected deviation (red – blue)/blue ch/MIP [ch] PMT ID Deviation

40 summary for physics applied new optimized HV (done) – #330555~ LUT update (done) – used run # update of constants of online monitor (done) mask for ch40 (done) ***PHYSICS ON*** starting, 2/12 12:30PM~ need to update DB need to update some plots in online monitor – please see Chirs mail /2/12


Download ppt "Log for RUN11 1. pedestal threshold scan Data for threshold scan were taken by Nihashi- kun. data: 5data points for theshold 5, 10, 15, 20, 25mV 2."

Similar presentations


Ads by Google